### v1

# Learning State-Invariant Representations of Objects from Image Collections with State, Pose, and Viewpoint Changes

Abstract
We add one more invariance — the state invariance — to the more commonly used other invariances for learning object representations for recognition and retrieval. By state invariance, we mean robust with respect to changes in the structural form of the object, such as when an umbrella is folded, or when an item of clothing is tossed on the floor. Since humans generally have no difficulty in recognizing objects despite such state changes, we are naturally faced with the question of whether it is possible to devise a neural architecture with similar abilities. To that end, we present a novel dataset, ObjectsWithStateChange, that captures state and pose variations in the object images recorded from arbitrary viewpoints. We believe that this dataset will facilitate research in fine-grained object recognition and retrieval of 3D objects that are capable of state changes. The goal of such research would be to train models capable of generating object embeddings that remain invariant to state changes while also staying invariant to transformations induced by changes in viewpoint, pose, illumination, etc. To demonstrate the usefulness of the ObjectsWithStateChange dataset, we also propose a curriculum learning strategy that uses the similarity relationships in the learned embedding space after each epoch to guide the training process. The model learns discriminative features by comparing visually similar objects within and across different categories, encouraging it to differentiate between objects that may be challenging to distinguish due to changes in their state. We believe that this strategy enhances the model’s ability to capture discriminative features for fine-grained tasks that may involve objects with state changes, leading to performance improvements on object-level tasks not only on the new dataset we present, but also on two other challenging multi-view datasets such as ModelNet40 and ObjectPI.
Index Terms:
Deep Metric Learning, Object Recognition and Retrieval, Pose-invariant and State-invariant Embeddings, Fine-grained Tasks, Multi-modal Dataset with Multi-view Images of Objects in Different States and Text Descriptions
1 Introduction
The fact that the appearance of an object depends significantly on its pose vis-a-vis that of the camera has received much attention in the research literature. However, there exists a more confounding cause for major changes in object appearance — the change in the state of the object. Consider, for example, objects like umbrellas that may be open or closed, books on a table that may also be open or closed, furniture items that are likely to be folded up when not in use, items of clothing and linen in different forms depending on how they are put away, etc. Fig. 1 shows some examples of such objects from our ObjectsWithStateChange dataset.
Figure 1: In addition to the pose, many commonly occuring objects also exhibit significant changes in their appearance when their state changes. This figure displays several examples of such objects in our ObjectsWithStateChange dataset.
We posit that with regard to the objects we enounter in our daily lives, it is just as likely as not that you will encounter objects whose appearance depends significantly on what state we find them in. That makes important the following question: Can modern computer vision algorithms effectively recognize objects despite the changes in their state? That is, in addition to achieving pose and viewpoint invariances, is it possible to also achieve invariance with respect to state changes? We believe that the work we present is a step in the direction of answering those questions.
TABLE I: Comparison of the new ObjectsWithStateChange (OWSC) dataset with the existing multi-view object classification datasets.
Datasets Year Type of Data # Catg. # Objects Real- world objects In-the-wild environment (Complex b/g) Fine- grained Arbitrary views State changes Train and Test split
Train Test
ModelNet40 [1] 2015 CAD 40 3183 800
FG3D [2] 2021 CAD 66 21575 3977
ETH-80 [3] 2003 RGB 8 80 80
RGB-D [4] 2011 RGB+Depth 51 300 300
MIRO [5] 2018 RGB 12 120 120
Toybox [6] 2018 Video 3 12 12
ObjectPI [7] 2019 RGB 25 382 98
ObjectNet [8] 2019 RGB 313 - -
RPC [9] 2022 RGB 17 200 200
MVP-N [10] 2022 RGB - 44 44
ObjectsWith
StateChange (Ours)
2024 RGB+Text 21 331 331
Obviously, any exploration of the answers to the questions posed above must begin with the creation of a dataset that captures not only pose and viewpoint appearance changes, but also state changes. To that end, a primary goal of this paper is the presentation of a new dataset we have created — the ObjectsWithStateChange dataset (hrta. OWSC). We believe that this dataset nicely augments the other well-known datasets — ModelNet40, ObjectPI, and FG3D — that are only meant to train and test for invariances with respect to pose and viewpoint. We believe that the OWSC dataset would facilitate creating image representations that would also be invariant to highly nonlinear shape deformations caused by state changes in addition to being invariant to changes in viewpoint, pose, illumination, etc.
To demonstrate the usefulness of the OWSC dataset, we also propose an extension to the dual-encoder model presented in [11] that learns object and category representations in separate embedding spaces simultaneously from multiple images of an object under different transformations. The model learns discriminative features by comparing images of a pair of objects that are randomly sampled from the same category. In order to accommodate the changes in object state, we extend [11] by proposing a curriculum learning strategy that uses the inter-object distances in the learned embedding space after each epoch to guide the sampling process. Specifically, this involves sampling images of similar objects from the same category (by retrieving neighboring objects from the same category, with and without state changes, in the learned embedding space) and from diverse categories (by dividing the learned embedding space into finer partitions as training progresses and sampling neighboring objects from the same partition). By doing so, we encourage the model to differentiate between harder-to-distinguish objects, enhancing its ability to capture discriminative features for fine-grained tasks in the presence of changes in state.
In summary, our main contributions are:
  • We introduce a new dataset called ObjectsWithStateChange that incorporates state changes in object appearance that are in addition to the more commonly used transformations related to changes in pose and viewpoint. We evaluate several state-of-the-art approaches on eight different category and object-level classification and retrieval tasks using this dataset.
  • We propose a new mining strategy to learn discriminative object-level representations more effectively that appears to work well even in the presence of state changes for the objects. This strategy leads to performance improvements on object-level tasks not only on the proposed ObjectsWithStateChange dataset but also on two other publicly available multi-view datasets, ObjectPI and ModelNet-40.
2 Related Work
(A) Multi-view Object Datasets: Prior work has introduced several multi-view object datasets for recognition and retrieval tasks. Datasets such as ModelNet-40 [1], and FG3D [2] comprise multi-view images of CAD models of synthetic objects. These datasets are intended for learning shape of objects and lack color and texture information, that we see in real-world objects. Datasets such as ETH-80 [3], W-RGBD [4], RPC [9], MIRO [12], MVP-N [10] capture real-world objects placed in a controlled environment with uniform illumination and a plain background while the ObjectPI [7] dataset captures objects in the wild with complex backgrounds. In all of these datasets, multi-view data is captured of objects placed in a particular pose from pre-defined viewpoints. ObjectNet [8] is a test dataset meant for evaluation of category recognition in a real-world setting with arbitrary object backgrounds, rotations, and viewpoints. Toybox [6] is a small video dataset capturing basic transformations such as rotation, translation, and zoom.
In contrast, our OWSC dataset captures complex state changes in addition to alterations in pose, lighting, background, and viewpoint, that induce significant changes in the appearance of the objects within the dataset. Notably, our dataset includes visually similar objects as shown in Fig. 2, making it well-suited for fine-grained tasks. A comparison of OWSC with other datasets is detailed in Table I. Furthermore, to aid in the learning of multi-modal representations, we provide text annotations describing the visual attributes of the objects.
(B) Deep Metric Learning Losses and Exemplar Mining:
Ranking Losses:
In deep metric learning, ranking losses [13, 14, 15] aim to optimize relative pairwise distances among exemplars, pulling similar samples closer together while pushing dissimilar ones apart. Selecting informative exemplars [16, 17, 18, 19, 20, 21, 22, 23] is crucial but adds computational costs and memory. To reduce the training complexity proxy-based methods [24] define a proxy embedding for each class and optimize the sample-to-proxy distances, which however does not capture the rich sample-to-sample relations that are particularly crucial for fine-grained tasks.
Curriculum Learning: As introduced in the seminal work [25], it organizes learning tasks by gradually increasing complexity and presenting easier examples before more challenging ones [26]. In the domain of metric learning, Sanakoyeu et al. [27, 28] partition both the data and the embedding space into smaller subsets and learn separate distance metrics within each of the non-overlapping subspaces. Meanwhile, Sarkar et al. [29, 30] proposed a hierarchical sampling strategy for retrieval tasks, initially sampling negatives from the same high-level category and then selecting harder negatives from the same fine-grained category.
Figure 2: This figure shows a few examples of visually similar objects from different categories in our dataset that makes it suitable for fine-grained recognition and retrieval tasks.
Pose-invariant Ranking Losses: Multi-view [31, 5, 32, 1] and pose-invariant methods [7] primarily focused on learning embeddings specific to categories, where the object-to-object variations are represented by variations in embedding vectors within the same embedding space. However, these approaches struggle to effectively distinguish between embeddings of different objects within the same category, leading to subpar performance on object-based tasks. To ameliorate this, PiRO [11] designed a dual-encoder model that explicitly decouples object and category embeddings. By training this model using pose-invariant ranking losses optimized in separate embedding spaces simultaneously, PiRO achieved improved object recognition and retrieval performance by effectively distinguishing between similar objects within the same category.
In our work, we evaluate the performance of these methods on our new dataset. Additionally, we introduce a modified training approach that employs the same losses as PiRO [11], but with a curriculum learning strategy that follows a different mechanism for sampling informative examples and conducting joint training, as detailed in Sec. 4.2.1. While PiRO randomly samples a pair of objects from the same category, we additionally sample similar objects from the same and other categories. As the training progresses, we sample objects that are harder to distinguish. This approach captures more discriminative object-identity representations to distinguish between visually similar objects, leading to performance improvements on object-level tasks, as shown later in Tables III, III.
(C) Transformer-based Architectures: Hybrid architectures, combining a CNN backbone with a transformer encoder [33], have been used for learning aggregated representations from image collections comprising different items [30, 29] and multi-view 3D shapes [34, 35] for classification and retrieval tasks. PiRO [11] introduced a dual-encoder to simultaneously learn pose-invariant category and object representations. While PiRO utilizes a single-head self-attention layer for each subspace to aggregate visual features from different views, we adopt a similar architecture with multiple attention layers. This modification allows us to effectively capture invariant visual features under more complex transformations beyond just pose changes.
3 The ObjectsWithStateChange Dataset
3.1 Dataset Design and Collection
Category and Object Selection: Objects are chosen such that they satisfy the following two conditions: (a) Objects have significant appearance changes under different state and pose changes (b) Objects are chosen such that different objects from the same category have very subtle differences for fine-grained object recognition and retrieval tasks as shown in Fig. 2. From 21 categories [^1] The categories are bags, books, bottles, bowls, clothes, cups, decorations, headphones, telephones, pillows, plants, plates, remotes, retail products, toys, ties, towels, trolley bags, tumblers, umbrellas, and vegetables., 331 objects that meet these criteria are selected. For each object, multiple images are collected based on the method specified next.
Dataset Collection: The OWSC dataset consists of images of 331 household objects captured using smartphone cameras. For each object, two sets of images are collected:
  • Transformation [^2] Note that ‘transformation’ in the context of our work refers to variations in state as well as viewpoint, pose, illumination, background, etc. split: For this split, images of each object in diverse states are captured from multiple arbitrary viewpoints. The objects are placed in multiple stable poses on different surfaces, featuring diverse backgrounds and varying lighting conditions. Detailed descriptions outlining the different transformations specific to each category are elaborated in the supplemental (ref. Table V).
  • Probe split: For this split, images of each object placed in an unseen state are captured from arbitrary viewpoints. The state of the objects, including the background and pose, differs from the images in the transformation split.
Due to differences in the physical attributes of the objects, they may undergo diverse types of transformations, resulting in a dataset with a variable number of images per object. Overall, the transformation split comprises 7900 images with an average of 24 images per object, and the probe split comprises 3428 images with approximately 10 images for each object. The transformation split is used for training, while the probe split comprising images of the same objects in an unseen state, is used for testing. Some examples are shown later in Fig. 7 and Figs. 8, 9 in the supplemental.
Dataset Annotation and Pre-processing: Each object is annotated with a category label, object label, and a text description describing the visual characteristics of the object to facilitate multi-modal learning. The images are captured from arbitrary viewpoints without recording specific pose or viewpoint information. No preprocessing is performed to preserve the natural background of the images, ensuring their suitability for in-the-wild recognition and retrieval.
3.2 Benchmarking on ObjectsWithStateChange
Single-image and Multi-image Invariant Recognition and Retrieval Tasks: We evaluate methods on eight invariant category and object-level recognition and retrieval tasks on our dataset. Single-image or multi-image category recognition. These tasks predict the category from a single image or a set of object images in an unseen state respectively. Single-image or multi-image object recognition. These tasks predict the object label from a single image or a set of object images in an unseen state respectively. Single-image or multi-image category retrieval. The goal of these tasks is to retrieve images of objects from the same category as the query object from a single image or multiple images respectively. Single-image or multi-image object retrieval. The goal of these tasks is to retrieve images of the same object-identity as the query object from a single image or multiple images respectively. Note that for the retrieval tasks, the query images (from the probe split) are in a different state as compared to the retrieved images (from the transformation split). The data splits will be publicly released at https://github.com/sarkar-rohan/ObjectsWithStateChange.
Evaluation metrics: Classification performance is measured in terms of accuracy and retrieval performance is measured using mean average precision (mAP).
Figure 3: An overview of the ranking losses in [11] for the object space (top) and the category space (bottom) that we use in our work.
Figure 3: An overview of the ranking losses in [11] for the object space (top) and the category space (bottom) that we use in our work.
Figure 4: An overview of the proposed training pipeline to learn invariant embeddings. Our curriculum learning strategy samples pairs of visually similar objects from the same and other categories to train the dual encoder jointly using the losses in Fig. 4.
4 Learning Invariant Embeddings
Sec. 4.1 provides a brief overview of various state-of-the-art pose-invariant methods. In our work, we evaluate their performance on our ObjectsWithStateChange (OWSC) dataset. Further, we extend the state-of-the-art method [11], as detailed in Sec. 4.2.
4.1 Background of Prior Pose-invariant Methods
The methods discussed in this section were originally designed to learn pose-invariant embeddings from multi-view images, but are adaptable to learn invariant embeddings from object image collections with diverse transformations from arbitrary views as they cluster multiple images of the same object without needing explicit pose or viewpoint information during training.
4.1.1 Single Embedding Space Methods:
Ho et al. [7] proposed three pose-invariant methods PI-CNN, PI-Proxy, and PI-TC that primarily focused on learning category-specific embeddings, where the object-to-object variations are represented by variations in embedding vectors within the same embedding space. Specifically, these methods clustered single-view embeddings close to the multi-view embedding of the same object, that is clustered close to a learned proxy embedding of its category. However, the embeddings of different objects within the same category are not explicitly separated, yielding poor performance on object-based tasks.
4.1.2 Dual Embedding Space Method:
To address this, PiRO [11] designed a dual-encoder model that explicitly decouples category and object embeddings. It comprises a shared CNN backbone, and two attention heads for learning category and object-identity representations in separate embedding spaces. The attention heads capture shared features that remain invariant across various transformations of the same object using self-attention. The encoder takes as input a set of $V$ unordered images of an object undergoing transformations and outputs $V$ single-image embeddings for each individual image, and an aggregated multi-image embedding for each embedding space.
The model is trained using three specialized margin-based losses for the object and category embedding spaces. In Fig. 4, we illustrate how the same losses in [11] can be used for learning object and category representations that are invariant to transformations and state changes. The losses are:
Object Loss: The loss $L_{piobj}$ identifies confusing instances between pairs of objects, shown as confusers in Fig. 4 (top). It pulls the confuser embeddings closer to the multi-image embedding of the same object, reducing intra-object distances (shown by green arrows), facilitating learning invariant features across transformations. Simultaneously, it increases inter-object distances by separating confuser embeddings and multi-image embeddings of different objects (shown by red arrows), aiding in learning discriminative features to distinguish between visually similar objects. This control over intra-object compactness and inter-object separability is achieved through margins $\alpha$ and $\beta$ respectively.
Category Losses: The loss $L_{picat}$ clusters embeddings of objects from the same category together in the category embedding space as shown by the green arrows in Fig. 4 (bottom). It ensures that the distances between the single-image and multi-image embeddings of an object, and between the multi-image embeddings of different object-identities within the same category do not exceed a margin $\theta$. Conversely, the loss $L_{cat}$ utilizes the large-margin softmax loss [36] to separate embeddings of objects from different categories by a margin $\gamma$ in the category embedding space, as shown by the red arrows in Fig. 4 (bottom).
The $L_{picat}$ and $L_{piobj}$ losses optimize the distances between the embeddings of a pair of objects. PiRO sampled these pairs of objects randomly from the same category and trained the dual-encoder to disentangle the category and object representations in the dual embedding spaces by learning common features between them in the category space and distinctive features between them in the object space simultaneously.
4.2 Extending PiRO using Curriculum Learning
We now extend the PiRO framework [11] as follows to deal with the challenges caused by state changes of objects that exhibit similar visual characteristics: First, while PiRO utilizes a single-headed self-attention layer in its dual encoder in order to aggregate visual features from different viewpoints, our new approach uses multiple attention layers. This adaptation enables us to effectively capture invariant visual features under complex transformations beyond mere changes in pose. Secondly, we introduce a modified training approach that uses the same losses as [11], detailed in Sec. 4.1.2, but with a new curriculum learning strategy based on a novel mechanism for mining and joint training involving the different loss functions. While PiRO randomly samples a pair of objects from the same category, now we additionally sample similar objects from the same and other categories. As the training progresses, we sample objects that are harder to distinguish (based on the inter-object distances in the embedding space) to effectively learn discriminative object-identity representations in order to distinguish between visually similar objects.
4.2.1 A New Curriculum Learning Training Strategy:
As shown in Fig. 4, the training begins by randomly sampling objects from the same category in the first epoch. Subsequently, at the start of each epoch, the learned multi-image object embeddings are used to identify similar objects from the same and other categories to guide the sampling process. Our method progressively samples objects that are closer in the learned embedding space as the training progresses.
Our training scheduler switches between the following three different sampling strategies to construct multi-image batches from pairs of objects for training:
(S1) Randomly sample objects from the same category:
Similar to PiRO [11], it allows sampling diverse objects from the same category that may be visually different (such as a pink handbag and a yellow lunch box), helping the encoder learn common features between them in the category space and discriminative features in the object space.
(S2) Randomly sample similar objects from the same category:
While the previous sampling strategy will include similar objects from the same category, the additional samples generated by this step increases the proportion of the visually similar pairs of objects from the same category in order to give greater weightage to such pairs. This allows the system to learn fine-grained differences between the similar-looking objects from the same category such as a pink handbag and a pink backpack from the category ‘bags’. At the start of the epoch, for each object in the dataset, a list of top-k neighboring objects in the learned object embedding space is retrieved, and, for pair formation, another object is randomly sampled from this list for training.
For (S1) and (S2), the dual encoder is jointly trained using the losses mentioned in Section 4.1.2, as follows
$L_{joint}=\frac{1}{|\mathcal{P}_{cat}|}\sum_{(x,y)\in\mathcal{P}_{cat}}L^{x}_{cat}+L^{y}_{cat}+L^{x,y}_{picat}+L^{x,y}_{piobj}$ (1)
where, $\mathcal{P}_{cat}$ is the set of all object pairs where each pair $(x,y)$ is sampled from the same category either randomly for (S1) or based on inter-object distances for (S2).
(S3) Randomly sample similar objects from any category:
At the start of the epoch, the object embedding space is partitioned into $f(N_{e})$ Voronoi cells using k-means, where $f$ is a function strictly increasing with the current epoch number $N_{e}$. Thus, the object embedding space is progressively divided into finer partitions as the training progresses. For each object in the dataset, another object is randomly sampled from the same partition. As training progresses, the pair of objects will be more difficult to separate in the object embedding space as the objects are going to be closer in finer partitions. This strategy allows sampling neighboring objects from any category that are visually similar in the object space helping the encoder learn discriminative features to distinguish between similar objects from any categories (such as a pink handbag and a pink T-shirt).
For (S3), the dual encoder is jointly trained using
$L_{joint}=\frac{1}{|\mathcal{P}_{part}|}\sum_{(x,y)\in\mathcal{P}_{part}}L^{x}_{cat}+L^{y}_{cat}+L^{x,y}_{piobj}$ (2)
where, $\mathcal{P}_{part}$ is the set of all object pairs where each pair comprises neighboring objects from the same partition in the learned object embedding space at the beginning of the epoch. Note that we do not use the $L^{x,y}_{picat}$ as the object pairs $(x,y)$ may be from different categories.
4.2.2 Training Details:
For a fair comparison with prior methods [7, 11], VGG-16 [37] is used as the CNN backbone. The last FC layers are modified to generate 2048-D embeddings and are initialized with random weights. Two-layer single-head self-attention layers are used for each embedding space with a dropout of 0.25. Training involves jointly optimizing the category and object-based losses, as described in Eqs. 1, 2. Margins $\alpha=0.25$ and $\beta=1.00$ are set for the object space, while margins $\theta=0.25$ and $\gamma=4.00$ are set for the category space, which are the same as [11]. $V$ is set to 12 images per object, and the images are resized to 224$\times$224 and normalized. We use the Adam optimizer with a learning rate of $5e^{-5}$. We train for 150 epochs and use the step scheduler that reduces the learning rate by half after every 30 epochs. For curriculum learning, the object embedding space is divided into $2N_{e}$ partitions, where $N_{e}$ is the current training epoch.
5 Experimental Results
5.1 Evaluation on ObjectsWithStateChange
TABLE II: Comparison of classification and retrieval performance for the category and object-based tasks against several state-of-the-art pose-invariant methods on the new ObjectsWithStateChange dataset. SV and MV denote either a single image or multiple images used at the time of inference respectively. The results indicate that our method outperforms prior methods on all tasks.
Method Embed. Space Classification (Accuracy %) Retrieval (mAP %)
Category Object Avg. Category Object Avg.
SV MV SV MV SV MV SV MV
PI-CNN Single 82.57 84.29 25.37 26.74 54.74 84.75 86.87 11.35 36.89 54.96
PI-Proxy Single 84.51 86.86 27.10 28.85 56.83 86.69 89.33 12.27 39.83 57.03
PI-TC Single 81.55 84.29 42.04 52.12 65.00 84.21 88.52 22.02 61.72 64.12
PiRO Dual 88.07 92.30 69.01 74.78 81.04 89.44 94.16 68.97 81.86 83.61
Ours Dual 89.20 93.66 76.31 85.35 86.13 90.59 95.00 78.00 89.46 88.26
We compare our method with prior pose-invariant methods (PI-CNN, PI-Proxy, and PI-TC) by Ho et al. [7] that learn pose-invariant embeddings in the single embedding space. We also compare against the current state-of-the-art, PiRO [11] that learns pose-invariant category and object embeddings in two separate spaces, resulting in improved performance on object-level tasks.
Figure 5: Comparison of the UMAP visualization of the embeddings of the 331 objects in our OWSC dataset learned using different methods. The embeddings for each object are shown using a unique color and pattern. The figures in the top row show the embeddings learned by the pose-invariant methods [7] that learn a single embedding (SE) space while the figures in the bottom row show the embeddings for methods learning a dual embedding (DE) space. It can be seen that the SE methods cannot separate objects from the same category. In comparison, the dual embedding (DE) methods can better separate the different objects despite the transformations. When comparing PiRO (bottom-left) and our proposed method (bottom-right), we see that the embeddings for different objects are better separated using our method.
We train all of the prior methods on our transformation split and test on the probe split. The results are presented in Table II and we can observe that our method outperforms all the prior methods. This is because our method can separate fine-grained objects under different transformations more effectively than prior methods as demonstrated using UMAP visualization of the embeddings in Fig. 5. For our method, we can observe that the embeddings of each object are better clustered and separated from other objects. This is primarily due to the curriculum learning strategy that not only separates objects from the same category but also separates visually similar objects from other categories. Also, in the next subsection, we observe that using multiple attention layers further improves performance as it helps capture invariant features under complex transformations and state changes. We provide detailed ablation of both the curriculum learning and the architectural changes in Section 5.2. We further observe that our curriculum learning approach improves performance on object-level tasks on two other publicly available multi-view datasets (ObjectPI [7] and ModelNet-40 [1]) in Table III.
TABLE III: Ablation study to understand the effect of sampling training examples. PiRO [11] randomly sampled a pair of objects from the same category, whereas our method additionally samples a pair of neighboring objects from the same and other categories. We observe that our curriculum learning strategy improves recognition and retrieval performance on object-level tasks as compared to the state-of-the-art [11].
(a) Results on single and multi-image invariant recognition and retrieval tasks on our ObjectsWithStateChange (OWSC) dataset.
Dataset Sampling Method Classification (Accuracy %) Retrieval (mAP %)
Category Object Avg. Category Object Avg.
SV MV SV MV SV MV SV MV
OWSC Same category 87.1 90.9 68.7 76.4 80.8 88.6 93.2 68.8 82.9 83.4
Curriculum (Ours) 86.9 89.4 76.6 83.4 84.1 88.6 92.4 78.0 88.0 86.8
Dataset Sampling Method Classification (Accuracy %) Retrieval (mAP %)
Category Object Avg. Category Object Avg.
SV MV SV MV SV MV SV MV
Object PI Same category 71.3 83.7 92.7 98.0 86.4 65.7 83.4 81.0 99.0 82.3
Curriculum (Ours) 70.3 85.7 96.2 99.0 87.8 64.9 82.4 86.3 99.3 83.2
Model Net40 Same category 85.2 88.9 93.7 96.9 91.2 79.7 86.1 84.0 98.2 87.0
Curriculum (Ours) 84.7 89.0 95.5 97.5 91.7 79.6 85.8 87.9 98.6 88.0
(a) Results on single and multi-image invariant recognition and retrieval tasks on our ObjectsWithStateChange (OWSC) dataset.
(b) Results on single and multi-view pose-invariant recognition and retrieval tasks on two publicly available multi-view datasets.
5.2 Ablation Studies
(A) Sampling Training Examples: PiRO uses multi-view images of a pair of objects randomly sampled from the same category, whereas our method additionally samples images of a pair of neighboring objects from the same category and other categories. We compare the performance of the two sampling strategies on the new ObjectsWithStateChange dataset on transformation-invariant recognition and retrieval tasks in Table III. We additionally also compare performance on pose-invariant recognition and retrieval tasks on two publicly available multi-view datasets (ObjectPI [7] and ModelNet-40 [1]) in Table III. For a fair comparison, the same architecture with a single-layer and single-head attention layer is used in both cases.
We observe that our strategy improves recognition and retrieval performance on object-based tasks. Specifically, in Table III, it improves single-view object recognition by 7.9% and single-view object retrieval by 9.2% over sampling objects randomly from the same category [11]. Similarly, in Table III, we observe that our sampling strategy improves single-view object recognition accuracy by 3.5% on ObjectPI and 1.8% on ModelNet, and single-view object retrieval mAP by 5.3% on ObjectPI and 3.9% on ModelNet. For multi-view object recognition and retrieval tasks, we see a significant improvement on our OWSC dataset and marginal improvements on ObjectPI and ModelNet-40 datasets.
However, we observe slight performance degradation on some of the category-based tasks. We conjecture that our curriculum learning approach focuses more on separating visually similar objects to improve performance on object-level tasks than category-level tasks. While these samples are harder to separate in the object embedding space, they are much easier to pull together in the category space as they have very similar attributes. Therefore, the samples that are informative for learning discriminative features between visually similar objects are not informative for learning common attributes between objects in the category space, thereby degrading performance on category tasks.
TABLE IV: Ablation study to understand the effect of the number of heads and self-attention layers in the architecture. Using two layers and a single head yields the best performance on all tasks.
Encoder Classification (Accuracy %) Retrieval (mAP %)
Category Object Avg. Category Object Avg.
nH nL SV MV SV MV SV MV SV MV
1 1 86.9 89.4 76.6 83.4 84.1 88.6 92.4 78.0 88.0 86.8
2 1 87.8 91.2 75.2 84.3 84.6 89.3 93.7 76.8 88.9 87.2
4 1 89.2 92.2 75.4 85.5 85.6 90.4 94.4 77.3 89.8 88.0
8 1 88.7 91.8 76.6 85.2 85.6 90.0 93.7 78.1 89.2 87.8
1 2 89.2 93.7 77.0 86.7 86.6 90.5 94.7 78.7 90.2 88.5
2 2 89.4 92.2 77.5 85.8 86.2 90.7 94.2 79.1 89.6 88.4
4 2 88.9 93.1 77.3 85.2 86.1 90.6 95.0 78.9 89.1 88.4
8 2 89.7 91.8 78.6 86.4 86.6 90.6 93.3 80.1 90.4 88.6
Figure 6: Variation of $\tau_{info}$ (top), and ratio $\rho$, and distances $d^{max}_{intra}$ and $d^{min}_{inter}$ (bottom) vs epochs during training on our OWSC dataset.
Figure 6: Variation of $\tau_{info}$ (top), and ratio $\rho$, and distances $d^{max}_{intra}$ and $d^{min}_{inter}$ (bottom) vs epochs during training on our OWSC dataset.
Figure 7: This figure (best viewed zoomed in) presents the results of single-image object retrieval from our dataset using our proposed method. The blue bounding box highlights the query image, while the green bounding box indicates the successful retrieval of an image of the correct object-identity. Our method demonstrates accurate retrieval of images of the same object under different transformations in (a) on the left side and effectively distinguishes the correct object from similar-looking ones as shown in consecutive rows in (b) on the right side.
(B) Architecture: As compared to [11] that used a single attention layer with a single head, we experimented with using multiple attention layers and multi-head attention to capture invariant visual features under complex transformations beyond just pose changes in the case of multi-view datasets. The results in Table IV indicate that using a single attention layer and multiple heads improves multi-image recognition and retrieval performance marginally. Using two attention layers with a single head yields the best performance. However, increasing the number of heads does not lead to any significant improvements for category-level tasks but yields marginal improvements on object-level tasks.
[…]
Similar to [11], we plot the maximum intra-class distance ($d^{max}_{intra}$ in green), the minimum inter-class distance between object-identity classes ($d^{min}_{inter}$ in blue), and the ratio $\rho=\frac{d^{min}_{inter}}{d^{max}_{intra}}$ (in red) during training in Fig. 7 (bottom). These metrics are used to monitor the compactness and separability of object-identity classes. Specifically, a lower $d^{max}_{intra}$, and higher $d^{min}_{inter}$ and $\rho$ values would indicate embeddings of the same object-identity class are well clustered and separated from embeddings of other object-identity classes. We observe that both sampling methods decrease $d^{max}_{intra}$ to the same extent. However, the curriculum learning strategy increases $d^{min}_{inter}$ and thereby increases $\rho$ to a greater extent than random sampling. Therefore, the object-identity classes are better separated (indicated by the blue lines in the bottom figure), and this results in improved performance on object-level tasks.
(D) Retrieval Results: The single-image object retrieval results on our dataset presented in Fig. 7, show that our method can accurately retrieve images of the same object under various transformations in (a) on the left side and also effectively distinguish the correct object from similar ones in (b) on the right side. Figs. 8, 9 in the supplemental show more single-image retrieval results of visually similar objects from each category in our dataset.
6 Conclusion
We present ObjectsWithStateChange, a novel dataset that, in addition to capturing the more commonly used transformations like those related to pose and viewpoint changes, also includes changes in the state of the objects. We evaluate several state-of-the-art methods across eight category-level and object-level classification and retrieval tasks on this dataset. Additionally, we introduce an example mining strategy to learn more discriminative object-level representations for fine-grained tasks. This approach not only boosts performance on our dataset but also on two other multi-view datasets available publicly.
Broader Impact and Future Work: Firstly, we introduce a novel dataset that captures complex state changes of objects that can be used for category-level and fine-grained object-level recognition and retrieval. This could benefit several real-world applications such as automatic checkout systems [38, 39, 9] and robotic systems that deal with object state changes upon interaction. Secondly, it can serve as a real-world test dataset for assessing the robustness of representations learned to state and pose changes, as we show in the supplemental in Appendix B. Lastly, we provide text annotations that can be used to learn invariant multi-modal representations[40, 41] and can be used for generalized zero-shot evaluation, which we aim to pursue in future work.
Appendix A State changes of different objects
A comprehensive list of all possible state changes and other transformations for objects from each category is provided in Table V. Depending on the physical characteristics of the specific object, all the transformations may not be possible for each object, and images are collected if such state changes are possible for the objects of interest.
Appendix B Cross-Dataset Evaluation for Assessing Robustness to State Changes and Other Transformations
Cross-dataset evaluation involves training a model on one dataset and then testing it on a different dataset. It is used to assess the generalization ability of models, particularly to evaluate robustness and reliability across different data sources or conditions that were not present in the training data. In this regard, our dataset can be used solely as an evaluation dataset, for testing the robustness of learned representations to state changes and other transformations in real-world scenarios.
To demonstrate this, we use the PiRO model [11] and our method trained on the ObjectPI [7] dataset and test it on the proposed ObjectsWithStateChange dataset and show the results in Table VI. Note that between the two datasets, there are 5 seen and 16 unseen categories and all unseen objects. Even for the seen categories, the object transformations and state changes are not present in the training data and the viewpoints are arbitrary. We observe that our curriculum learning slightly improves performance on object-level tasks.
TABLE VI: Results of cross-dataset performance evaluation. Model is trained on ObjectPI and tested on the ObjectsWithStateChange dataset with 5 seen and 16 unseen categories and all unseen objects. Even for the seen categories, the object transformations and state changes are unseen, and the viewpoints are arbitrary.
Sampling Method Classification (Accuracy %) Retrieval (mAP %)
Category Object Avg. Category Object Avg.
SV MV SV MV SV MV SV MV
Same category 65.3 64.7 43.7 45.9 54.9 30.2 42.7 17.9 55.1 36.5
Curriculum (Ours) 64.2 63.4 46.5 49.6 55.9 29.6 42.2 20.8 58.6 37.8
Here we compared the performance of two methods trained on the same ObjectPI dataset and tested on our proposed dataset. In a similar manner, it is also possible to compare the performance of models trained on different datasets to assess how robust and generalizable models trained on different datasets are.
Appendix C Qualitative Fine-grained Retrieval Results
As mentioned earlier in Section 3, our dataset comprises visually similar objects from each of the 21 categories. In this section, we show some qualitative single-image retrieval results of visually similar objects from each category in our dataset in Figs. 8 and 9.
In these figures, the results are presented in two columns that show the results for similar-looking objects from the same category in each row. Firstly, this illustrates that our dataset has several similar-looking objects from each category with very subtle differences in appearance and hence can facilitate research in fine-grained object retrieval tasks. Secondly, the results indicate that our algorithm is able to retrieve images of the same object-identity with high mAP despite having similar-looking objects from each category and different state changes and other transformations from arbitrary viewpoints.
Appendix D Discussion
Broader Impact and Future Research: Firstly, we believe that to the best of our knowledge, we are the first to introduce a dataset that captures complex state changes of objects that can be used for category-level and fine-grained object-level recognition and retrieval tasks. This could benefit several real-world applications such as automatic checkout systems [38, 39, 9], robotic systems, that deal with object state changes upon interaction. Secondly, this dataset can be used for cross-dataset evaluation of methods trained on different datasets and tested for their robustness to transformations captured in our ObjectsWithStateChange dataset. Lastly, we provide text annotations describing the visual characteristics of objects that can be used for training multi-modal models such as [40, 41]. This can be used for adapting such models on object image collections for learning transformation-invariant multi-modal representations and generalized zero-shot evaluation of algorithms, which we aim to pursue in future.
Challenges: The following aspects make our dataset challenging: (a) significant appearance changes due to state changes and other transformations such as arbitrary pose, viewpoint, lighting, and background changes (b) presence of several similar-looking objects with fine-grained differences which makes recognition and retrieval of the correct object-identity challenging especially when there are state changes (c) Additionally, there is class imbalance as some categories such as headphones, telephones, and umbrellas have less number of objects as compared to clothes, decorations, retail objects as objects in the latter categories are more easily available in households. Also, retail objects and decorations have a lot more images than other categories such as cups, bottles, and plates, as many transformations due to pose and state changes are possible in the former than the latter categories. Due to these reasons, the dataset is imbalanced both with respect to the number of objects per category and the number of images per object, which makes it even more challenging. (d) images of the objects are captured in the wild with cluttered and complex backgrounds which facilitates evaluation of algorithms in real-world scenarios.